# Kilohertz Gravitational Waves from Binary Neutron Star Mergers: Full Spectrum Analyses and High-density Constraints on Neutron Star Matter

### v1

@@ II.3 EoS constraints from quasi-universal relations
Subsequently, we use a sample of two millions EoSs from Godzieba et al. (2021) generated under minimal assumptions, i.e. general relativity, causality, the maximum NS mass above the maximum NS mass from pulsar observations, ${M_{\rm max}^{\rm TOV}}>1.97{\rm M_{\odot}}$ Antoniadis et al. (2013), and consistent with the upper bound on the reduced tidal parameter $\tilde{\Lambda}$ from GW170817 Abbott et al. (2017a, 2019b, 2018b). We reweigh the EoS sample with the posteriors of the masses, the reduced tidal parameter, ${\rho_{\rm max}^{\rm TOV}}$ and ${R_{\rm max}^{\rm TOV}}$, employing emcee sampler Foreman-Mackey et al. (2013). This allows us to place constraints not only on the NS mass-radius diagram, but also to determine the 90% credibility region for the tidal deformability curve $\Lambda(M)$ and on the pressure-density relation $p(\rho)$. Specifically, from the EoS sampling we obtain a distribution of EoSs that most closely resemble the one used for the injection. For each sampled EoS, we compute the relations between the mass and the radius of the star and between the tidal polarizability parameter and the mass, yielding a 90% confidence interval (CI) on the $M(R)$ and the $\Lambda(M)$ curve, respectively. Using the same EoS distribution we solve the Tolmann-Oppenheimer-Volkoff (TOV) equations to determine the corresponding pressure-density relation, from which we derive the 90% confidence region for the $p(\rho)$ curve.
The maximum mass of non-rotating NSs, ${M_{\rm max}^{\rm TOV}}$, is a crucial parameter for constraining the mass-radius diagram and consequently the EoS. Additionally, it is also essential for the determination of the probability of prompt collapse. Specifically, PC is expected to occur when the total mass of the system exceeds a threshold mass, which is related through phenomenological relations to ${M_{\rm max}^{\rm TOV}}$ Hotokezaka et al. (2011); Bauswein et al. (2013):
@@
In our PC analyses, we consider equal-mass binaries and determine the $90\%$ CI of the maximum mass ${M_{\rm max}^{\rm TOV}}$ and the corresponding radius ${R_{\rm max}^{\rm TOV}}$ of the NS from the mass-radius diagram and insert them in Eq. (11). We then calculate the probability of PC as
@@ III Results
III.1 Inspiral-merger
Figure 3: Relative precision on chirp mass (top) and mass ratio (bottom) with respect to the different initial frequencies for a non-spinning BNS with $\mathcal{M}=1.1976917{\rm M_{\odot}}$, $q=1.5$, $\tilde{\Lambda}=488$, at different luminosity distances. A lower initial frequency or a closer source leads to improved precision in both the chirp mass and the mass ratio estimations.
@@ III.1 Inspiral-merger
Table 1: Summary of the PEs with TEOBResumSPA for a BNS with $\mathcal{M}=1.1976917{\rm M_{\odot}}$, $q=1.5$, $\tilde{\Lambda}=488$, $\delta\tilde{\Lambda}=-60$, $\chi_{\rm eff}=0$, $\iota=0$. The first three columns report the injected luminosity distance, the initial frequency of the analyses and the SNR of the injected data; the last seven columns report the recovered values, with the median of the posterior distributions and the $90\%$ CIs.
@@
We generate artificial data in the frequency domain, using the triangular, triple-interferometer ET detector Hild et al. (2008); Punturo et al. (2010b); Maggiore et al. (2020), with ET-D sensitivity Hild et al. (2011); Hild (2012) and a sampling rate of $4096~{}{\rm Hz}$. The sky location corresponds to the optimal location for the ET detector $\{\alpha=2.64,\delta=0.762\}$ and the binaries are oriented with $\iota=0$ and $\psi=0$. The injections are performed in zero-noise, including only the dominant (2,2) mode. The injected strain is analyzed in the frequency domain from $f_{0}$ to $2048~{}{\rm Hz}$, where $f_{0}\in[2,5,10,20]~{}{\rm Hz}$. We set the priors for chirp mass $\mathcal{M}\in[0.8,1.5]{\rm M_{\odot}}$, for mass ratio $q\in[1,2]$ and for the tidal polarizability parameters $\Lambda_{1,2}\in[0,3000]$. The spins are aligned with the orbital angular momentum, with an isotropic prior $\chi_{1,2}\in[-0.05,0.05]$ for the $z-$component. Finally, for the luminosity distance we employ volumetric priors $D_{L}\in[10,150]~{}{\rm Mpc}$, except when the injected luminosity distance is $300~{}{\rm Mpc}$, where we use $D_{L}\in[200,400]~{}{\rm Mpc}$. We marginalize over reference phase.
@@ III.1.1 Source parameters
where $x_{m}$ is the median value and $x_{95\%}$ ($x_{5\%}$) corresponds to the $95$th ($5$th) percentile of the posterior distribution of the parameter $x$. The injected GW signals correspond to non-spinning BNS with $\mathcal{M}=1.1976917{\rm M_{\odot}}$, $q=1.5$ and $\tilde{\Lambda}=488$. We vary the luminosity distance in $[40,80,120,300]~{}{\rm Mpc}$. The precision on both $\mathcal{M}$ and $q$ increases as the distance of the source decreases, since the SNR is higher. Similarly, keeping a fixed luminosity distance, the precision increases as the initial frequency $f_{0}$ of the analysis decreases, e.g. at $D_{L}=300~{}{\rm Mpc}$, decreasing the frequency from $20~{}{\rm Hz}$ to $5~{}{\rm Hz}$, the precision on the measurement of both chirp mass and mass ratio improves of about one order of magnitude.
@@
with respect to the SNR, for eight different binaries. As the SNR increases, that is when the luminosity distance decreases, the width of the distribution shrinks, meaning that the measurement of $\tilde{\Lambda}$ is more precise. This is mostly due to the improvement in the chirp mass and mass ratio discussed above. We run the same test fixing the luminosity distance and varying the initial frequencies of the analyses. The results are presented in Fig. 5. We obtain a similar result as before: as the SNR increases, namely as the initial frequency decreases, it is possible to have a better estimation on the value of $\tilde{\Lambda}$, e.g. of about a factor two, when the initial frequency changes from 20 to $5~{}{\rm Hz}$. Reducing the luminosity distance from $120~{}{\rm Mpc}$ to $40~{}{\rm Mpc}$, instead, improves the precision on the reduced tidal parameters by a factor between two and three. At the same time, we notice that having a larger luminosity distance, i.e. 80 Mpc, but a smaller initial frequency, i.e. 5 Hz, enhances the precision in the measurement of all the parameters with respect to a closer source ($D_{L}=40~{}{\rm Mpc}$) at $f_{0}=20~{}{\rm Hz}$, even if the SNR is higher in the latter case. Thus, decreasing the initial frequency is an effective strategy to achieve better constraints in the estimation of the binary parameters.
Table 1 presents the recovered median values with the $90\%$ confidence regions for the inferred parameters. In addition to the consideration above about the chirp mass and the mass ratio, as the initial frequency decreases from 20 Hz to 2 Hz, the precision on the estimation of the reduced tidal parameter increases of about a factor two and the precision on the effective spin of one order of magnitude. The precision of the luminosity distance and of the inclination, instead, is comparable for the different cases and the inferred values are not always compatible with the injected ones. This is due to the known degeneracy between distance and inclination, which arises when considering waveforms constructed only with the $(2,\pm 2)$ modes Cutler and Flanagan (1994). Despite this, our results of the precision of the luminosity distance are consistent with those reported in Singh and Bulik (2022).
In Table 1 and in Table LABEL:tab:IM2 of Appendix C we also report the recovered values of the asymmetric tidal parameter $\delta\tilde{\Lambda}$. The estimation on this parameter carries substantial uncertainties, in some case larger than 100%. The trend in the improvement of its precision follows the same behavior discussed above for the reduced tidal parameter. We notice that, when the SNR is sufficiently high, i.e. SNR1500greater-than-or-equivalent-toabsent1500$\gtrsim$ 1500≳ 1500 for $q>1.2$, the recovered values of $\delta\tilde{\Lambda}$ deviate significantly from zero, indicating a measurable asymmetry in the tidal polarizability parameters. Furthermore, the joint estimation of $\delta\tilde{\Lambda}$ and $\tilde{\Lambda}$ leads to constraints on the individual tidal polarizability parameters. In few cases, when the mass ratio is $q\gtrsim 1.289$ and the SNR has the highest value, i.e. $f_{0}=$ 5Hz and $D_{L}=$ 40Mpc, we notice that both the posterior distributions of $\Lambda_{1}$ and $\Lambda_{2}$ are not compatible with zero, enabling us to confidently identify the source as a BNS merger. However, in all the other cases, the posterior distribution of $\Lambda_{1}$ remains compatible with zero, preventing us from discarding the BH-NS hypothesis.
@@ III.1.2 Comparison of ET configurations
In Figure 5 we compare two different configurations of ET, the triangular one Hild et al. (2011); Hild (2012), described in Sec. II.2 and a two-L configuration Branchesi et al. (2023), where the first detector is located in the Meuse-Rhine region and the second one in Sardinia, with the same sensitivity as the triangular one Hild et al. (2011); Hild (2012) and a relative orientation between the two detectors of $45^{\circ}$. We find that the posterior distributions of $\tilde{\Lambda}$ for the two configurations are comparable. Our results are in line with Puecher et al. (2023b).
@@
Branchesi et al. (2023); Begnoni et al. (2025) and Puecher et al. (2023b) compared various ET geometries using a Fisher matrix approach in the former cases and a Bayesian PE in the latter. They all concluded that an improvement in the precision on the inference of the source parameters comes primarily from the arms length, while changing shape or alignment of the detectors plays a minimal role in the accuracy increase. Additionally, Branchesi et al. (2023) found that the estimation of the source parameters is comparable for the two configurations we considered, while the triangular performs better in the estimation of the inclination angle. Regarding the angular localization, their results indicate that the two-L is better compared to the triangular one. This improvement is attributed to the ability of the two-L setup to partially triangulate, thanks to its longer baseline. These findings are compatible with our results.
@@ III.2 Inspiral-merger-postmerger
Table 2: Summary of the PEs with TEOBResumSPA for IM and TEOBResumSPA_NRPMw for IMPM case. The first five columns report the injected parameters, included the EoS used; the central six columns report the recovered values, with the median of the posterior distributions and the $90\%$ CIs and the last column displays the BFs, with statistical uncertainties as in Skilling (2006).
@@
This section is dedicated to the analysis of IMPM injection-recovery PEs. We employed the complete TEOBResumSPA_NRPMw template for the injections of GWs from different binaries, i.e. with different EoSs. As in the IM-only case, the detector used is ET, with triangular shape and ET-D sensitivity Hild et al. (2011); Hild (2012). The sky location is the same as in III.1 and the luminosity distance is $40~{}{\rm Mpc}$. The sampling rate is $16384~{}{\rm Hz}$ and the analyses are performed in the frequency range $f\in[20,8192]~{}\rm{Hz}$. The priors for chirp mass are $\mathcal{M}\in[0.8,2]\,{\rm M_{\odot}}$, for mass ratio $q\in[1,4]$, for the tidal polarizability parameters $\Lambda_{1,2}\in[0,4000]$ and for the luminosity distance are volumetric $D_{L}\in[10,1000]~{}{\rm Mpc}$. We perform the inference under the assumption of non-spinning binaries and we keep all the other extrinsic parameters fixed to the injected values. We add the recalibration parameters $\boldsymbol{\delta}_{\rm fit}$ to account for intrinsic errors of the calibrated formulae (see Paper I).
@@
Figure 8 presents five illustrative recovered spectra, in which the injected waveforms are included within the $90\%$ CIs of the reconstructed spectra. Particularly, both the inspiral part and the frequency of the peak of the PM, $f_{2}$, are well recovered with tight error bars for all the binaries, i.e. it is possible to obtain constraints on $f_{2}$ for different EoSs with an uncertainty smaller than 8%. The high-frequency features and the region between the merger and the peak of the PM are consistent with the injection, but, in these cases, the $90\%$ CI are broader, indicating lower precision on the detection compared to the inspiral segment. This reduced precision arises because the strain amplitude is about two orders of magnitude smaller and the detector sensitivity is lower in the high-frequency region.
@@
where $p(d|H)$ is the evidence. In the upper part of the Table the BFs are larger than zero, meaning that the IMPM model is preferred over the IM model. All the BF values are on the order of ten, whereas the associated uncertainties, estimated with the criterion introduced in Skilling (2006), are on the order $10^{-1}$. This indicates a high degree of statistical confidence, supporting a clear detection of the presence of the PM. In the bottom part of the Table, instead, we test the ability of our pipeline to recover PC. The relative discussion is postponed to Sec. III.4.

### v2

@@ II.3 EoS constraints from quasiuniversal relations
Subsequently, we use a sample of two millions EoSs from Godzieba et al. (2021) generated under minimal assumptions, i.e. general relativity, causality, the maximum NS mass above the maximum NS mass from pulsar observations, ${M_{\rm max}^{\rm TOV}}>1.97{\rm M_{\odot}}$ Antoniadis et al. (2013), and consistent with the upper bound on the reduced tidal parameter $\tilde{\Lambda}$ from GW170817 Abbott et al. (2017a, 2019b, 2018). We reweigh the EoS sample with the posteriors of the masses, the reduced tidal parameter, $\rho_{\rm max}^{\rm TOV}$} and $R_{\rm max}^{\rm TOV}$}, employing emcee sampler Foreman-Mackey et al. (2013). This allows us to place constraints not only on the NS mass-radius diagram, but also to determine the 90% credibility region for the tidal deformability curve $\Lambda(M)$ and on the pressure-density relation $p(\rho)$. Specifically, from the EoS sampling we obtain a distribution of EoSs that most closely resemble the one used for the injection. For each sampled EoS, we compute the relations between the mass and the radius of the star and between the tidal polarizability parameter and the mass, yielding a 90% confidence level (CL) on the $M(R)$ and the $\Lambda(M)$ curve, respectively. Using the same EoS distribution we solve the Tolmann-Oppenheimer-Volkoff (TOV) equations to determine the corresponding pressure-density relation, from which we derive the 90% confidence region for the $p(\rho)$ curve.
The maximum mass of nonrotating NSs, $M_{\rm max}^{\rm TOV}$}, is a crucial parameter for constraining the mass-radius diagram and consequently the EoS. Additionally, it is also essential for the determination of the probability of prompt collapse. Specifically, PC is expected to occur when the total mass of the system exceeds a threshold mass, which is related through phenomenological relations to $M_{\rm max}^{\rm TOV}$} Hotokezaka et al. (2011); Bauswein et al. (2013)
@@
In our PC analyses, we consider equal-mass binaries and determine the $90\%$ CL of the maximum mass $M_{\rm max}^{\rm TOV}$} and the corresponding radius $R_{\rm max}^{\rm TOV}$} of the NS from the mass-radius diagram and insert them in Eq. (11). We then calculate the probability of PC as
@@ III Results
III.1 Inspiral merger
Figure 3: Relative precision on chirp mass (top) and mass ratio (bottom) with respect to the different initial frequencies for a nonspinning BNS with $\mathcal{M}=1.1976917{\rm M_{\odot}}$, $q=1.5$, $\tilde{\Lambda}=488$, at different luminosity distances. A lower initial frequency or a closer source leads to improved precision in both the chirp mass and the mass ratio estimations.
@@ III.1 Inspiral merger
Table 1: Summary of the PEs with TEOBResumSPA for a BNS with $\mathcal{M}=1.1976917{\rm M_{\odot}}$, $q=1.5$, $\tilde{\Lambda}=488$, $\delta\tilde{\Lambda}=-60$, $\chi_{\rm eff}=0$, $\iota=0$. The first three columns report the injected luminosity distance, the initial frequency of the analyses and the SNR of the injected data; the last seven columns report the recovered values, with the median of the posterior distributions and the $90\%$ CLs.
@@
We generate artificial data in the frequency domain, using the triangular, triple-interferometer ET detector Hild et al. (2008); Punturo et al. (2010b); Maggiore et al. (2020), with ET-D sensitivity Hild et al. (2011); Hild (2012) a sampling rate of $4096~{\rm Hz}$, and a duration of $128~{\rm s}$. We tested also longer durations to ensure that the length of the signal does not have any effects on the results. For example, with a signal length of 512 s and initial frequency $5~{\rm Hz}$, the posteriors are compatible within 90% confidence regions to the ones shown below, since both injection and recovery are in the frequency domain. The sky location corresponds to the optimal location for the ET detector $\{\alpha=2.64,\delta=0.762\}$ and the binaries are oriented with $\iota=0$ and $\psi=0$. The injections are performed in zero noise, including only the dominant (2,2) mode. The injected strain is analyzed in the frequency domain from $f_{0}$ to $2048~{\rm Hz}$, where $f_{0}\in[2,5,10,20]~{\rm Hz}$. We set the priors for chirp mass $\mathcal{M}\in[0.8,1.5]{\rm M_{\odot}}$, for mass ratio $q\in[1,2]$ and for the tidal polarizability parameters $\Lambda_{1,2}\in[0,3000]$. The spins are aligned with the orbital angular momentum, with an isotropic prior $\chi_{1,2}\in[-0.05,0.05]$ for the $z$-component. Finally, for the luminosity distance we employ volumetric priors $D_{L}\in[10,150]~{\rm Mpc}$, except when the injected luminosity distance is $300~{\rm Mpc}$, where we use $D_{L}\in[200,400]~{\rm Mpc}$. We used a uniform distribution for all the priors and we marginalize over reference phase. In principle it is possible to use priors for the tidal polarizability parameters that consider $\Lambda_{1}\leq\Lambda_{2}$, as shown in Fig. 1 of Abbott et al. (2018). We estimate the impact of this possible choice by removing samples from our posterior that do not satisfy this condition, which is equivalent to imposing this condition as a prior and repeating the analysis. We notice a mildly impact in the results of the inference of the polarizability tidal parameters, obtaining a smaller 90% credible intervals for $\tilde{\Lambda}$ and $\delta\tilde{\Lambda}$. For instance, for the injection discussed in Table 1, the 90% credible interval for $\tilde{\Lambda}$ changes from [481,499] to [481,498] at $D_{L}=40$Mpc and $f_{0}=2$Hz and from [430,610] to [450,610] at $D_{L}=300$Mpc and $f_{0}=20$Hz when imposing $\Lambda_{1}<\Lambda_{2}$. The area of the posterior distributions of the tidal polarizability parameters of Figs. 4 and 5 shrinks as well, the difference, though, is really slight and not even present in all the cases. Therefore, we employed agnostic priors on $\Lambda$ for the results shown.
@@ III.1.1 Source parameters
where $x_{m}$ is the median value and $x_{95\%}$ ($x_{5\%}$) corresponds to the $95$th (55th) percentile of the posterior distribution of the parameter $x$. The injected GW signals correspond to nonspinning BNS with $\mathcal{M}=1.1976917{\rm M_{\odot}}$, $q=1.5$, $\tilde{\Lambda}=488$ and $\delta\tilde{\Lambda}=-60$. These values are compatible with GW170817 event Abbott et al. (2019a). We vary the luminosity distance in $40,80,120,300]~{\rm Mpc}$. The precision on both $\mathcal{M}$ and $q$ increases as the distance of the source decreases, since the SNR is higher. Similarly, keeping a fixed luminosity distance, the precision increases as the initial frequency $f_{0}$ of the analysis decreases, e.g. at $D_{L}=300~{\rm Mpc}$, decreasing the frequency from $20~{\rm Hz}$ to $5~{\rm Hz}$, the precision on the measurement of both chirp mass and mass ratio improves of about one order of magnitude.
@@
with respect to the SNR, for eight different binaries. As the SNR increases, that is when the luminosity distance decreases, the width of the distribution shrinks, meaning that the measurement of $\tilde{\Lambda}$ is more precise. This is mostly due to the improvement in the chirp mass and mass ratio discussed above. We run the same test fixing the luminosity distance and varying the initial frequencies of the analyses. The results are presented in Fig. 5. We obtain a similar result as before: as the SNR increases, namely as the initial frequency decreases, it is possible to have a better estimation on the value of $\tilde{\Lambda}$, e.g. of about a factor two, when the initial frequency changes from 20 to $5~{\rm Hz}$. Reducing the luminosity distance from $120~{\rm Mpc}$ to $40~{\rm Mpc}$, instead, improves the precision on the reduced tidal parameters by a factor between two and three. At the same time, we notice that having a larger luminosity distance, i.e., 80 Mpc, but a smaller initial frequency, i.e., 5 Hz, enhances the precision in the measurement of all the parameters with respect to a closer source ($D_{L}=40~{\rm Mpc}$) at $f_{0}=20~{\rm Hz}$, even if the SNR is higher in the latter case. Thus, decreasing the initial frequency is an effective strategy to achieve better constraints in the estimation of the binary parameters.
Table 1 presents the recovered median values with the $90\%$ confidence regions for the inferred parameters. In addition to the consideration above about the chirp mass and the mass ratio, as the initial frequency decreases from 20 Hz to 2 Hz, the precision on the estimation of the reduced tidal parameter increases of about a factor two and the precision on the effective spin of one order of magnitude. The precision of the luminosity distance and of the inclination, instead, is comparable for the different cases and the inferred values are not always compatible with the injected ones. This is due to the known degeneracy between distance and inclination, which arises when considering waveforms constructed only with the $(2,\pm 2)$ modes Cutler and Flanagan (1994) Despite this, our results of the precision of the luminosity distance are consistent with those reported in Singh and Bulik (2022).
[…]
@@ III.1.2 Comparison of ET configurations
In Fig. 5 we compare two different configurations of ET, the triangular one Hild et al. (2011); Hild (2012), described in Sec. II.2 and a two-L configuration Branchesi et al. (2023), where the first detector is located in the Meuse-Rhine region and the second one in Sardinia, with the same sensitivity as the triangular one Hild et al. (2011); Hild (2012) and a relative orientation between the two detectors of $45^{\circ}$. We find that the posterior distributions of $\tilde{\Lambda}$ for the two configurations are comparable. Our results are in line with Puecher et al. (2023b).
@@
Branchesi et al. (2023); Begnoni et al. (2025), and Puecher et al. (2023b) compared various ET geometries using a Fisher matrix approach in the two former cases and a Bayesian PE in the latter two. They all concluded that an improvement in the precision on the inference of the source parameters comes primarily from the arms length, while changing shape or alignment of the detectors plays a minimal role in the accuracy increase. Additionally, Branchesi et al. (2023) found that the estimation of the source parameters is comparable for the two configurations we considered, while the triangular performs better in the estimation of the inclination angle. Regarding the angular localization, their results indicate that the two-L is better compared to the triangular one. This improvement is attributed to the ability of the two-L setup to partially triangulate, thanks to its longer baseline. These findings are compatible with our results.
@@ III.2 Inspiral-merger-postmerger
Table 2: Summary of the PEs with TEOBResumSPA for IM and TEOBResumSPA_NRPMw for IMPM case. The first five columns report the injected parameters, included the EoS used; the central six columns report the recovered values, with the median of the posterior distributions and the $90\%$ CLs and the last column displays the BFs, with statistical uncertainties as in Skilling (2006).
@@
This section is dedicated to the analysis of IMPM injection-recovery PEs. We employed the complete TEOBResumSPA_NRPMw template for the injections of GWs from different binaries, i.e. with different EoSs. As in the IM-only case, the detector used is ET, with triangular shape and ET-D sensitivity Hild et al. (2011); Hild (2012). The sky location is the same as in III.1 and the luminosity distance is $40~{\rm Mpc}$. The sampling rate is $16384~{\rm Hz}$ and the signal length $128~\rm{s}$, the analyses are performed in the frequency range $f\in[20,8192]~\rm{Hz}$. The priors for chirp mass are $\mathcal{M}\in[0.8,2]\,{\rm M_{\odot}}$, for mass ratio $q\in[1,4]$, for the tidal polarizability parameters $\Lambda_{1,2}\in[0,4000]$ and for the luminosity distance are volumetric $D_{L}\in[10,1000]~{\rm Mpc}$. We perform the inference under the assumption of nonspinning binaries and we keep all the other extrinsic parameters fixed to the injected values. We add the recalibration parameters $\boldsymbol{\delta}_{\rm fit}$ to account for intrinsic errors of the calibrated formulas (see paper I).
@@
Figure 8 presents five illustrative recovered spectra, in which the injected waveforms are included within the $90\%$ CLs of the reconstructed spectra. Particularly, both the inspiral part and the frequency of the peak of the PM, $f_{2}$, are well recovered with tight error bars for all the binaries, i.e. it is possible to obtain constraints on $f_{2}$ for different EoSs with an uncertainty smaller than 8%. The high-frequency features and the region between the merger and the peak of the PM are consistent with the injection, but, in these cases, the $90\%$ CL are broader, indicating lower precision on the detection compared to the inspiral segment. This reduced precision arises because the strain amplitude is about two orders of magnitude smaller and the detector sensitivity is lower in the high-frequency region.
@@
[…]